OpenStudy (anonymous):

what wrong with this equation ? 2(n+1.5)=4.5 2n=6 n=3 please help

OpenStudy (shadowfiend):

So, let's start at the beginning. $$ 2(n + 1.5) = 4.5 $$ You can distribute the 2: $$ \begin{align} 2n + 2 \cdot (1.5) &= 4.5\\ 2n + 3 &= 4.5\\ \end{align} $$ Can you see how that does not lead to \(2n = 6\), and therefore your final answer isn't right?